OpenStudy (anonymous):

How do you do this the format is y-y1=m(X-X1) The problem: (-6,3), m= -1/2 I need help please help me.. :)

OpenStudy (anonymous):

Please i need help.

OpenStudy (anonymous):

You'd plug in the numbers into the formula. Y-3=-1/2(X+6) The -3 is Y1, and -6 is X1. Because X1 is -6 and the formula is X-X1, the -6 becomes positive. Distribute and get Y by itself in order to get your finished equation.

OpenStudy (anonymous):

Thank you. What wuld be the steps for that?

OpenStudy (anonymous):

Do you need help solving from that point?

OpenStudy (anonymous):

i just need the steps for leadin up to the answer thats it.

OpenStudy (anonymous):

After you plug in to get Y-3=-1/2(X+6), distribute the 1/2 to the x and 6. You should then have Y-3=-1/2X-3. Then get rid of the three on the Y side of the equation by canceling it out. Add that three to the other side. You should be finished then.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

No problem :)